OpenStudy (anonymous):

write the percent as a fraction or a mixed number in simplest from 12 5/9% help someone please!!!!

OpenStudy (paki):

is it (12*5/9%) or what.... @asiemariah

OpenStudy (anonymous):

no its 12 and 5/9

OpenStudy (paki):

12 is alone and 5/9 is alone....?

OpenStudy (anonymous):

no its all together

OpenStudy (paki):

so mean it is 12*5/9.... if is it (12*5/9%) then it will be equal to (113/9) / 100 = 113/900 = 0.12%

OpenStudy (anonymous):

I need to know what it is as a mixed number or a fraction it's already in a percent

OpenStudy (anonymous):

its a mixed number already but the mixed number is a percent if that makes any since

OpenStudy (paki):

yeah it is a mixed number and is equal to 12.555%

OpenStudy (anonymous):

im about to show you a picture just one second

OpenStudy (paki):

yeah, i have seen the picture... my answer is rite...